what is the dimensions of a rectangle with a perimeter of 70ft and length to width ratio of 4 to 5?
Answer by stanbon(75887)

what is the dimensions of a rectangle with a perimeter of 70ft and length to width ratio of 4 to 5?
----
4:5 is the same as 4x:5x where length = 4x and width = 5x
-------------------
Equation:
P = 2(length + width)
70 = 2(4x+5x)
35 = 9x
x = (35/9)
---
length = 4x = 4(35/9) = 140/9 = 15.56 ft
---
width = 5x = 5(35/9) = 175/9 = 19.44 ft.
=============================================
Cheers,
Stan H.